the length of a rectangle is 10 m more than its breadth. if the perimeter of rectangle is 80 m, find its dimensions.

Answer by Timnewman(323)

Let L be the lenght of thr rectangle,and w be the width of the rectangle.
from your question,
L=w+10---(1)
and,
2(L+w)=80
L+w=40----(2)
put equ 1 in 2
w+10+w=40
2w+10=40
2w=40-10
2w=30
2w/2 = 30/2
w=15m
substitute for w in equ (1)
L=w+10
L=15+10
L=25m
The lenght nd width of the rectangle is 15m nd 25m respectively.